# Fusion of Multiscale Features Via Centralized Sparse-attention Network for EEG Decoding

### v1

@@ 6.5 Hierarchical Guidance of Structure
Inspired by [55], we compared the performance of the main–auxiliary structure from EEG-CSANet with a hierarchical structure, as is shown in Figure 8. In the hierarchical design, the features extracted after the DW-Spa-Conv module in the first branch serve as the Query for the second branch’s MSCA, whose output is then passed sequentially as the Query to the third and fourth branches. This layer-by-layer propagation enables each branch to attend not only to its own representations but also to inherit information from the preceding branch, thereby achieving cross-branch multi-level feature fusion. As is shown in Table 11, compared with main–auxiliary structure, the Hierarchical structure yields slightly lower classification performance, though the difference is not statistically significant according to t-tests. To present the best-performing model, we ultimately adopted the main–auxiliary structure.
@@
Table 11: Comparison between Main–auxiliary Guidance and Hierarchical Guidance.

### v2

@@ 6.5 Hierarchical Guidance of Structure
Inspired by [liu2025vieeg], we compared the performance of the main–auxiliary structure from EEG-CSANet with a hierarchical structure, as is shown in Figure 8. In the hierarchical design, the features extracted after the DW-Spa-Conv module in the first branch serve as the Query for the second branch’s MSCA, whose output is then passed sequentially as the Query to the third and fourth branches. This layer-by-layer propagation enables each branch to attend not only to its own representations but also to inherit information from the preceding branch, thereby achieving cross-branch multi-level feature fusion. As is shown in Table 11, compared with main–auxiliary structure, the Hierarchical structure yields slightly lower classification performance, though the difference is not statistically significant according to t-tests. To present the best-performing model, we ultimately adopted the main–auxiliary structure.
@@
Table 11: Comparison between Main-auxiliary Guidance and Hierarchical Guidance.